You deposit $6000 in an account earning 4% interest compounded quarterly. How much will you have in the account in 5 years?
Answer by mananth(16946)

You deposit $6000 in an account earning 4% interest compounded quarterly. How much will you have in the account in 5 years?

Principal P = 6000
Amount= A
years=t 5.00
compounded 4 times a year n
Rate = 4.00 0.04 0.01
Amount = P*((n+r)/n)^n*t

Amount = = 6000 *( 1 + 0.01 )^ 4 * 5.00
Amount = 6000 *( 1 + 0.01 )^ 20
6000 *( 1.01 )^ 20.00
Amount = 7321.14